Given the functions: ATC= x-6+(30/x) and AR =10+(15/x), where x represents output, find TC, TR, and the output level when net revenue is zero.

1
Expert's answer
2021-02-18T15:36:30-0500

"ATC=\frac{TC}{Q}"

"AR=\frac{TR}{Q}"

Q=x

"TC=ATC\times Q"

"TR=ATR\times Q"

"TC=(x-6 + (\frac{30}{ x}))\times x=x^2-6x+30"

"TR= (10 + (\frac{15}{ x}))\times x=10x+15"

TR=TC=0

"(x^2-6x+30)-(10x+15)=0"

"x^2-16x+15=0"

x=1, x=15

x=15

TR=150+15=165

TC=225-90+30=165

x=1

TR=10+15=25

TC=1-6+30=25